sixteen men can completed a work in 15 days. 24 children can do the same in 20 days. in how many days will 8 men and 8 children together completed the same work.
Answer by stanbon(75887)

sixteen men can completed a work in 15 days. 24 children can do the same in 20 days. in how many days will 8 men and 8 children together completed the same work.
---------
16m * 15d = 1 job
16m = (1/15) job/day
8m = (1/30) job/day (this is the rate for 8 men)
-------------------------
24c * 20d = 1 job
24c = (1/20)job/day
8c = (1/60)job/day (this is the rate for 8 children)
---------------------
Let number of days for 8 men and 8 children be "x":
EQUATION:
x(1/30 + 1/60) = 1 job
x(3/60)(job/day) = 1 job
x = 20 days (this is time job would take 8 men and 8 children)
=====================
Cheers
Stan H.